# Poincaré polynomials of moduli spaces of one-dimensional sheaves on the projective plane

### v1

@@ Abstract.
For $S=\mathbb{P}^{2}$, our proof also introduces a new approach to computing the Poincaré polynomials via Gromov-Witten invariants of local $\mathbb{P}^{2}$ and a local elliptic curve. Specifically, we compute the Poincaré polynomials of $M_{d}$ with degrees $d\leq 16$ and derive a closed formula for the leading Betti numbers $b_{i}(M_{d})$ with $d\geq 6$ and $i\leq 4d-22$.
@@ 1.1. Background
The moduli space $M_{\beta,\chi}$ of of one-dimensional Gieseker semistable sheaves $\mathcal{F}$ on a del Pezzo surface $S$, with $[{\operatorname{Supp}}(\mathcal{F})]=\beta$ and Euler characteristic $\chi$, has attracted significant attention, as seen in works such as [LeP93, Y14, Y15, Y18, Y23, Y24, CC15, CC16, CvGKT, Bou22, Bou23, MS23, PS23, KPS23, LMP24, PSSZ24]. In this paper, we study the Poincaré polynomial of the intersection cohomology of $M_{\beta,\chi}$.
@@
to denote the Poincaré polynomial of $M_{\beta}$. Note that the odd Betti numbers of $M_{\beta}$ vanish by [M09]. The Poincaré polynomails $P_{\beta}\left(y\right)$ were conjectured to have the following divisibility property:
@@ 1.2. Results overview
which follows directly from an explicit formula (see Theorem 3.1 for more details) for stationary Gromov-Witten invariants of local elliptic curves. This formula is established by extending the topological quantum field theory (TQFT) formalism of [BP08] to include stationary insertions.
@@
There is a quite parallel story in [BW23]. Let $D$ be a smooth rational ample divisor in a smooth projective surface $X$. It was shown in [BW23, Theorem 1.7] that the generating series of maximal contact Gromov-Witten invariants of the pair $(X,D)$ (with fixed curve class $\beta$) can be divided by the (shifted) Poincaré polynomial of projective space $\mathbb{P}^{D\cdot\beta-1}$. Using the quiver/Gromov-Witten correspondence [Bou20, Bou21], this is equivalent to the fact that the (intersection) Poincaré polynomial of certain quiver moduli space can be divided by the Poincaré polynomial of $\mathbb{P}^{D\cdot\beta-1}$.
In contrast with with the case of $M_{\beta}$, there is a geometric explanation for the divisibility of Poincaré polynomial in the quiver case. Actually, it was shown in [BW23, Theorem 1.11] that such a quiver moduli space has a small resolution such that after the small resolution, it become a projective bundle.
@@ 1.3. Explicit calculations
Next, we specialize $S=\mathbb{P}^{2}$ and use $\Omega^{\mathbb{P}^{2}}_{d}$ to denote the shifted Poincaré polynomial of the moduli space $M_{d}$ of Gieseker semistable sheaves on $\mathbb{P}^{2}$ supported on curves with degree $d\in\mathbb{Z}_{>0}$. The equality (1) also introduces a noval approach to computing the Poincaré polynomials $\Omega^{\mathbb{P}^{2}}_{d}$ via Gromov-Witten invariants of local $\mathbb{P}^{2}$ and stationary Gromov-Witten invariants of a local elliptic curve. In Section 6, we give a detailed computation of $\Omega^{\mathbb{P}^{2}}_{d}$ for $d\leq 6$ using this new method, and provide a list of $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ for $d\leq 10$ where
@@
Actually, we have calculated $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ for $d\leq 16$ using Gopakumar-Vafa invariants of local $\mathbb{P}^{2}$ in [HKR]. Since the coefficients of $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ grow rapidly, we limit the list to $d\leq 10$ in this paper. A complete list of $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ for $d\leq 16$ is available at
@@
Here and below, we use $|_{y^{\leq m}}$ to denote truncation, retaining only the terms where the power of $y$ is at most $m$. Theorem 1.6 will be proven in Section 5, and the main ideas can be summarized as follows:
A $y$-degree analysis (see Propositions 5.1 and 5.3) on the RHS of (1) reveals that only two rooted trees contributes to the leading Betti numbers. Theorem 1.6 then follows from a closed formula for the leading Gopakumar-Vafa invariants, derived by the first author and Zhou in [GZ15] using the topological vertex formula (see Theorem 5.4 for more details).
@@
In [Y23, Y24], Yuan showed that the Betti numbers $b_{i}(M_{d})$ can be computed from the Betti numbers of Hilbert schemes of points on $\mathbb{P}^{2}$ when $i\leq 2d$. Yuan’s results can be reformulated as
@@
where $f(y)$ is independent of $d$, with the leading expansion
$f(y)=-2-2\,y+8\,{y}^{3}+17\,{y}^{4}+30\,{y}^{5}+42\,{y}^{6}+57\,{y}^{7}+69\,{y}^{8}+84\,{y}^{9}+96\,{y}^{10}+O({y}^{11}).$
A potential resolution of the above conjecture may require a more refined analysis of the Gopakumar-Vafa invariants of local $\mathbb{P}^{2}$. We leave this for future exploration.
@@ 1.4. Organization of the paper
This paper is organized as follows. In Section 2, we give a new way to package the all-genus local/relative correspondence. In Section 3, we give an explicit formula for Gromov-Witten invariants of local elliptic curves with stationary insertions. In Section 4, we give a proof of Theorem 1.3, and we give a careful $y$-degree analysis for the contributions arising in the graph sum formula established in Section 2. In Section 5, we give a proof of Theorem 1.6. In Section 6, we give a detailed computation of the shifted Poincaré polynomial $\Omega^{\mathbb{P}^{2}}_{d}$ for $d\leq 6$ and provide a list of $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ up to $d\leq 10$.
@@ 1.5. Acknowledgment
We would like to thank Pierrick Bousseau, Jinwon Choi, Michel van Garrel, Rahul Pandharipande, Junliang Shen for helpful discussions and comments. The first author is supported by by the National Key R&D Program of China (No. 2023YFA1009802) and NSFC 12225101. The second author is supported by the National Key R&D Program of China (No. 2022YFA1006200) and SwissMAP. The second author is especially grateful to Rahul Pandharipande for hosting an academic visit at ETH Zurich, which greatly facilitated the completion of this work.
@@ 2.2. A graph sum formula
    $t:V_{T}\longrightarrow\{0,1\}$, where we use $V_{T}$ to dentoe the set of vertices.
@@ 4. Applications of the all-genus local/relative correspondence
$=\sum_{k|\beta}\frac{-1}{k(y^{\frac{k}{2}}-y^{-\frac{k}{2}})^{2}}\sum_{g\geq 0}n_{g,\frac{\beta}{k}}(-1)^{g}(y^{\frac{k}{2}}-y^{-\frac{k}{2}})^{2g}$
@@ 4.2. Further analysis of the graph sum formula
$G_{0,\boldsymbol{\beta}}^{E}=\begin{cases}1,\text{ if }s=1,\\ 0,\text{ otherwise}\end{cases}$
@@
(15) $\begin{split}1+\sum_{\rho}\prod_{\square\in\rho}(-1)^{K_{S}^{2}}q^{[E]}y^{c(\square)K_{S}^{2}}\frac{I(qy^{-c(\square)K_{S}},p_{\beta_{i}},y)^{2}}{I(qy^{-(c(\square)+1)K_{S}},p_{\beta_{i}},y)I(qy^{-(c(\square)-1)K_{S}},p_{\beta_{i}},y)}&\\ =H(q,p_{\beta_{i}},y)\end{split}$
@@ 5.1. Degree analysis
$\mathcal{Z}_{d}$ $=\frac{(1-y)^{2}}{\prod_{k>0}(1-y^{k})^{3}}\cdot\left(\binom{d+2}{2}-3\sum_{i\geq 1}\frac{y^{i}}{(1-y^{i})^{2}}\right)$
@@
=\ = $\mathcal{Z}_{d}-\mathcal{Z}_{d-3}+3\left(-y^{d-1}\cdot\mathcal{Z}^{\prime}_{d}+y^{d-4}\cdot\mathcal{Z}^{\prime}_{d-3}-y^{d-4}(1+y+y^{2})^{2}\cdot\mathcal{Z}_{d-4}\right)+O({y^{2d-10}})$
@@ 6. Numerical results
$\frac{3d(-1)^{d^{2}+1}\Omega_{d}^{\mathbb{P}^{2}}\left(y^{\frac{1}{2}}\right)}{{\frac{\left(y^{\frac{1}{2}}\right)^{3d}-\left(y^{\frac{1}{2}}\right)^{-3d}}{y^{\frac{1}{2}}-y^{-\frac{1}{2}}}}}-\sum_{g\geq 0}n_{g,d}(-1)^{g}(y^{\frac{1}{2}}-y^{-\frac{1}{2}})^{2g}=0$
@@
$\Omega_{1}^{\mathbb{P}^{2}}$ $={\frac{y^{\frac{3}{2}}-y^{-\frac{3}{2}}}{y^{\frac{1}{2}}-y^{-\frac{1}{2}}}}$
$\Omega_{2}^{\mathbb{P}^{2}}$ $={\frac{y^{3}-y^{-3}}{y^{\frac{1}{2}}-y^{-\frac{1}{2}}}}$
@@
$\frac{-1}{(y^{\frac{1}{2}}-y^{-\frac{1}{2}})^{2}}\left(\frac{3d(-1)^{d^{2}+1}\Omega_{d}^{\mathbb{P}^{2}}\left(y^{\frac{1}{2}}\right)}{{\frac{\left(y^{\frac{1}{2}}\right)^{3d}-\left(y^{\frac{1}{2}}\right)^{-3d}}{y^{\frac{1}{2}}-y^{-\frac{1}{2}}}}}-\sum_{g\geq 0}n_{g,d}(-1)^{g}(y^{\frac{1}{2}}-y^{-\frac{1}{2}})^{2g}\right)=\sum_{T\in\mathcal{R}T_{d},\atop t^{-1}(0)\neq\emptyset}{\operatorname{Cont}}_{T}$
@@
${\operatorname{Cont}}_{T}=-F^{E,\operatorname{tw}}_{1,\emptyset}=1$
@@
$\Omega_{3}^{\mathbb{P}^{2}}=\frac{1}{y}(1+y+y^{2}){\frac{y^{\frac{9}{2}}-y^{-\frac{9}{2}}}{y^{\frac{1}{2}}-y^{-\frac{1}{2}}}}\$
@@
$\frac{-1}{(y^{\frac{1}{2}}-y^{-\frac{1}{2}})^{2}}\left(\frac{3d(-1)^{d^{2}+1}\Omega_{d}^{\mathbb{P}^{2}}\left(y^{\frac{1}{2}}\right)}{{\frac{\left(y^{\frac{1}{2}}\right)^{3d}-\left(y^{\frac{1}{2}}\right)^{-3d}}{y^{\frac{1}{2}}-y^{-\frac{1}{2}}}}}-\sum_{g\geq 0}n_{g,d}(-1)^{g}(y^{\frac{1}{2}}-y^{-\frac{1}{2}})^{2g}\right)=\sum_{T\in\mathcal{R}T_{d},\atop t^{-1}(0)\neq\emptyset}{\operatorname{Cont}}_{T}$
@@
$-\frac{2\sin(\frac{3\hbar}{2})}{3}F^{E,\operatorname{tw}}_{1,\{1\}}\mathcal{W}_{1}=-\frac{y^{3/2}-y^{-3/2}}{3i}\cdot\frac{3(y^{3/2}-y^{-3/2})}{i}\cdot\frac{-3}{(y^{1/2}-y^{-1/2})^{2}}=-3\left(\frac{y^{3/2}-y^{-3/2}}{y^{1/2}-y^{-1/2}}\right)^{2}$
@@
$n_{0,4}=-192,\quad n_{1,4}=231,\quad n_{2,4}=-102,\quad n_{3,4}=15$
@@
$\Omega_{4}^{\mathbb{P}^{2}}=\frac{1}{y^{3}}(1+y+4y^{2}+4y^{3}+4y^{4}+y^{5}+y^{6}){\frac{y^{6}-y^{-6}}{y^{\frac{1}{2}}-y^{-\frac{1}{2}}}}\$
@@
$\frac{-1}{(y^{\frac{1}{2}}-y^{-\frac{1}{2}})^{2}}\left(\frac{3d(-1)^{d^{2}+1}\Omega_{d}^{\mathbb{P}^{2}}\left(y^{\frac{1}{2}}\right)}{{\frac{\left(y^{\frac{1}{2}}\right)^{3d}-\left(y^{\frac{1}{2}}\right)^{-3d}}{y^{\frac{1}{2}}-y^{-\frac{1}{2}}}}}-\sum_{g\geq 0}n_{g,d}(-1)^{g}(y^{\frac{1}{2}}-y^{-\frac{1}{2}})^{2g}\right)=\sum_{T\in\mathcal{R}T_{d},\atop t^{-1}(0)\neq\emptyset}{\operatorname{Cont}}_{T}$
@@
${\operatorname{Cont}}_{T_{1}}=-\frac{2\sin(3\hbar)}{6}F^{E,\operatorname{tw}}_{1,\{2\}}\mathcal{W}_{2}=-\frac{3(y^{3}-y^{-3})^{2}}{2(y-y^{-1})^{2}}+\frac{6(y^{3}-y^{-3})^{2}}{(y^{1/2}-y^{-1/2})^{2}}$
@@
${\operatorname{Cont}}_{T_{2}}=-\frac{1}{2}\left(\frac{2\sin(3\hbar/2)}{3}\right)^{2}F^{E,\operatorname{tw}}_{1,\{1,1\}}(\mathcal{W}_{1})^{2}=\frac{9(y^{3/2}-y^{-3/2})^{4}}{2(y^{1/2}-y^{-1/2})^{4}}$
@@
${\operatorname{Cont}}_{T_{3}}$ $=\left(y^{9/2}-y^{-9/2}\right)^{2}\left(-\frac{1}{(y^{3/2}-y^{-3/2})^{2}}-\frac{10y+10y^{-1}+7}{(y^{1/2}-y^{-1/2})^{2}}\right)$
@@
This method allows us to determine $\Omega_{d}^{\mathbb{P}^{2}}$ once the Gopakumar-Vafa invariants $n_{g,d^{\prime}}$ for $d^{\prime}\leq d$ are known. In the following table, we provide a list of $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ for $d\leq 10$, where $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ is given by (2). Actually, we have computed $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ for $d\leq 16$ using Gopakumar-Vafa invariants of local $\mathbb{P}^{2}$ in [HKR]. Howerver, as the coefficients of $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ grow rapidly, we limit the list to $d\leq 10$ in this paper. A complete list of $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ for $d\leq 16$ is available at:

### v2

@@ Abstract.
For $S=\mathbb{P}^{2}$, our proof also introduces a novel approach to computing the Poincaré polynomials using Gromov-Witten invariants of local $\mathbb{P}^{2}$ and a local elliptic curve. Specifically, we compute the Poincaré polynomials of $M_{d}$ with degrees $d\leq 16$ and derive a closed formula for the leading Betti numbers $b_{i}(M_{d})$ with $d\geq 6$ and $i\leq 4d-22$. We also propose a conjectural formula for the leading Betti numbers $b_{i}(M_{d})$ with $d\geq 4$ and $i\leq 6d-20$. In the Appendix (by M. Moreira), a more general conjecture concerning the higher range Betti numbers of $M_{d}$ is presented, along with another conjecture that involves refinements from the perverse/Chern filtration.
@@ 1.1. Background
The moduli space $M_{\beta,\chi}$ of of one-dimensional Gieseker semistable sheaves $\mathcal{F}$ on a del Pezzo surface $S$, with $[{\operatorname{Supp}}(\mathcal{F})]=\beta$ and Euler characteristic $\chi$, has attracted significant attention, as seen in works such as [LeP93, Y14, Y15, Y18, Y23, Y24, CC15, CC16, CvGKT, Bou22, Bou23, MS23, PS23, KPS23, LMP24, klmp, PSSZ24]. In this paper, we study the Poincaré polynomial of the intersection cohomology of $M_{\beta,\chi}$.
@@
to denote the Poincaré polynomial of $M_{\beta}$. Note that the odd Betti numbers of $M_{\beta}$ vanish by [B95, M09]. The Poincaré polynomials $P_{\beta}\left(y\right)$ were conjectured to have the following divisibility property:
@@ 1.2. Results overview
which follows from an explicit formula (see Theorem 3.1 for more details) for stationary Gromov-Witten invariants of local elliptic curves. This formula is established by extending the topological quantum field theory (TQFT) formalism of [BP08] to include stationary insertions.
@@
There is a quit parallel story in [BW23]. Let $D$ be a smooth rational ample divisor in a smooth projective surface $X$. It was shown in [BW23, Theorem 1.7] that the generating series of maximal contact Gromov-Witten invariants of the pair $(X,D)$ (with fixed curve class $\beta$) can be divided by the (shifted) Poincaré polynomial of projective space $\mathbb{P}^{D\cdot\beta-1}$. Using the quiver/Gromov-Witten correspondence [Bou20, Bou21], this is equivalent to the fact that the (intersection) Poincaré polynomial of certain quiver moduli space can be divided by the Poincaré polynomial of $\mathbb{P}^{D\cdot\beta-1}$.
In contrast with with the case of $M_{\beta}$, there is a geometric explanation for the divisibility of Poincaré polynomial in the quiver case. Actually, it was shown in [BW23, Theorem 1.11] that such a quiver moduli space has a small resolution such that after the small resolution, it becomes a projective bundle.
@@ 1.3. Explicit calculations
Next, we specialize $S=\mathbb{P}^{2}$ and use $\Omega^{\mathbb{P}^{2}}_{d}$ to denote the shifted Poincaré polynomial of the moduli space $M_{d}$ of Gieseker semistable sheaves on $\mathbb{P}^{2}$ supported on curves with degree $d\in\mathbb{Z}_{>0}$. The equality (1) also introduces a novel approach to computing the Poincaré polynomials $\Omega^{\mathbb{P}^{2}}_{d}$ using Gromov-Witten invariants of local $\mathbb{P}^{2}$ and stationary Gromov-Witten invariants of a local elliptic curve. In Section 6, we give a detailed computation of $\Omega^{\mathbb{P}^{2}}_{d}$ for $d\leq 6$ using this new method, and provide a list of $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ for $d\leq 10$ where
@@
In fact, we have calculated $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ for $d\leq 16$ using Gopakumar-Vafa invariants of local $\mathbb{P}^{2}$ in [HKR]. Since the coefficients of $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ grow rapidly, we limit the list to $d\leq 10$ in this paper. A complete list of $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ for $d\leq 16$ is available at
@@
Throughout, $|_{y^{\leq m}}$ denotes truncation (keeping terms with $y$-powers $\leq m$). A geometric interpretation of the above formula in terms of the cohomology rings of $M_{d,\chi}$ is discussed in Appendix A.1. Theorem 1.6, proven in Section 5, can be summarized as follows:
A detailed $y$-degree analysis (see Propositions 5.1 and 5.3) on the RHS of (1) reveals that only two rooted trees contribute to the leading Betti numbers. Theorem 1.6 then follows from a closed formula for the leading Gopakumar-Vafa invariants, derived by the first author and Zhou in [GZ15] using the topological vertex formula (see Theorem 5.4 for more details).
@@
In [Y23, Y24], Yuan showed that the Betti numbers $b_{i}(M_{d})$ can be computed from the Betti numbers of Hilbert schemes of points in $\mathbb{P}^{2}$ when $i\leq 2d$. Yuan’s results can be reformulated as
@@
where
$f(y)=\frac{(1+y+y^{2})(-2+2y+4y^{2}+2y^{3}+y^{4}+2y^{5})}{(1-y)(1-y^{2})}.$
A potential resolution of the above conjecture may require a more refined analysis of Gopakumar-Vafa invariants of local $\mathbb{P}^{2}$. We leave this for future exploration.
Remark 1.10.
Conjecture 1.9 was initially formulated when $f(y)$ was known only up to order 10. The precise closed-form expression for $f(y)$ was proposed by M. Moreira based on our preliminary coefficients and certain geometric considerations. This further inspired Moreira to propose a more general conjecture on higher-range Betti numbers of $M_{d}$ (Conjecture A.4), as well as another conjecture (Conjecture A.5) involving refinements coming from the perverse/Chern filtration, both presented in Appendix A.
@@ 1.4. Organization of the paper
This paper is organized as follows. In Section 2, we give a new way to package the all-genus local/relative correspondence. In Section 3, we give an explicit formula for Gromov-Witten invariants of local elliptic curves with stationary insertions. In Section 4, we give a proof of Theorem 1.3, and we give a detailed $y$-degree analysis for the contributions arising in the graph sum formula established in Section 2. In Section 5, we give a proof of Theorem 1.6. In Section 6, we give a detailed calculation of the shifted Poincaré polynomial $\Omega^{\mathbb{P}^{2}}_{d}$ for $d\leq 6$ and provide a list of $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ up to $d\leq 10$. In Appendix A, a more general conjecture concerning the higher range Betti numbers of $M_{d}$ is presented, along with another conjecture that involves refinements from the perverse/Chern filtration.
@@ 1.5. Acknowledgment
We would like to thank Pierrick Bousseau, Jinwon Choi, Michel van Garrel, Rahul Pandharipande, Junliang Shen for helpful discussions and comments. Special thanks go to Miguel Moreira for contributing the Appendix and for his help in formulating Conjecture 1.9.
The first author is supported by the National Key R&D Program of China (No. 2023YFA1009802) and NSFC 12225101. The second author is supported by the National Key R&D Program of China (No. 2022YFA1006200) and SwissMAP. The second author is especially grateful to Rahul Pandharipande for hosting an academic visit at ETH Zurich, which greatly facilitated the completion of this work.
@@ 2.2. A graph sum formula
    $t:V_{T}\longrightarrow\{0,1\}$, where we use $V_{T}$ to denote the set of vertices.
@@ 4. Applications of the all-genus local/relative correspondence
$=\sum_{k|\beta}\frac{-1}{k(y^{\frac{k}{2}}-y^{-\frac{k}{2}})^{2}}\sum_{g\geq 0}n_{g,\frac{\beta}{k}}(-1)^{g}(y^{\frac{k}{2}}-y^{-\frac{k}{2}})^{2g}.$
@@ 4.2. Further analysis of the graph sum formula
$G_{0,\boldsymbol{\beta}}^{E}=\begin{cases}1,\text{ if }s=1,\\ 0,\text{ otherwise,}\end{cases}$
@@
(15) $\begin{split}1+\sum_{\rho}\prod_{\square\in\rho}(-1)^{K_{S}^{2}}q^{[E]}y^{c(\square)K_{S}^{2}}\frac{I(qy^{-c(\square)K_{S}},p_{\beta_{i}},y)^{2}}{I(qy^{-(c(\square)+1)K_{S}},p_{\beta_{i}},y)I(qy^{-(c(\square)-1)K_{S}},p_{\beta_{i}},y)}&\\ =H(q,p_{\beta_{i}},y).\end{split}$
@@ 5.1. Degree analysis
$\mathcal{Z}_{d}$ $=\frac{(1-y)^{2}}{\prod_{k>0}(1-y^{k})^{3}}\cdot\left(\binom{d+2}{2}-3\sum_{i\geq 1}\frac{y^{i}}{(1-y^{i})^{2}}\right),$
@@
=\ = $\mathcal{Z}_{d}-\mathcal{Z}_{d-3}+3\left(-y^{d-1}\cdot\mathcal{Z}^{\prime}_{d}+y^{d-4}\cdot\mathcal{Z}^{\prime}_{d-3}-y^{d-4}(1+y+y^{2})^{2}\cdot\mathcal{Z}_{d-4}\right)+O({y^{2d-10}}).$
@@ 6. Numerical results
$\frac{3d(-1)^{d^{2}+1}\Omega_{d}^{\mathbb{P}^{2}}\left(y^{\frac{1}{2}}\right)}{{\frac{\left(y^{\frac{1}{2}}\right)^{3d}-\left(y^{\frac{1}{2}}\right)^{-3d}}{y^{\frac{1}{2}}-y^{-\frac{1}{2}}}}}-\sum_{g\geq 0}n_{g,d}(-1)^{g}(y^{\frac{1}{2}}-y^{-\frac{1}{2}})^{2g}=0.$
@@
$\Omega_{1}^{\mathbb{P}^{2}}$ $={\frac{y^{\frac{3}{2}}-y^{-\frac{3}{2}}}{y^{\frac{1}{2}}-y^{-\frac{1}{2}}}}\,,$
$\Omega_{2}^{\mathbb{P}^{2}}$ $={\frac{y^{3}-y^{-3}}{y^{\frac{1}{2}}-y^{-\frac{1}{2}}}}\,.$
@@
$\frac{-1}{(y^{\frac{1}{2}}-y^{-\frac{1}{2}})^{2}}\left(\frac{3d(-1)^{d^{2}+1}\Omega_{d}^{\mathbb{P}^{2}}\left(y^{\frac{1}{2}}\right)}{{\frac{\left(y^{\frac{1}{2}}\right)^{3d}-\left(y^{\frac{1}{2}}\right)^{-3d}}{y^{\frac{1}{2}}-y^{-\frac{1}{2}}}}}-\sum_{g\geq 0}n_{g,d}(-1)^{g}(y^{\frac{1}{2}}-y^{-\frac{1}{2}})^{2g}\right)=\sum_{T\in\mathcal{R}T_{d},\atop t^{-1}(0)\neq\emptyset}{\operatorname{Cont}}_{T}\,.$
@@
${\operatorname{Cont}}_{T}=-F^{E,\operatorname{tw}}_{1,\emptyset}=1.$
@@
$\Omega_{3}^{\mathbb{P}^{2}}=\frac{1}{y}(1+y+y^{2}){\frac{y^{\frac{9}{2}}-y^{-\frac{9}{2}}}{y^{\frac{1}{2}}-y^{-\frac{1}{2}}}}\,.$
@@
$\frac{-1}{(y^{\frac{1}{2}}-y^{-\frac{1}{2}})^{2}}\left(\frac{3d(-1)^{d^{2}+1}\Omega_{d}^{\mathbb{P}^{2}}\left(y^{\frac{1}{2}}\right)}{{\frac{\left(y^{\frac{1}{2}}\right)^{3d}-\left(y^{\frac{1}{2}}\right)^{-3d}}{y^{\frac{1}{2}}-y^{-\frac{1}{2}}}}}-\sum_{g\geq 0}n_{g,d}(-1)^{g}(y^{\frac{1}{2}}-y^{-\frac{1}{2}})^{2g}\right)=\sum_{T\in\mathcal{R}T_{d},\atop t^{-1}(0)\neq\emptyset}{\operatorname{Cont}}_{T}\,.$
@@
$-\frac{2\sin(\frac{3\hbar}{2})}{3}F^{E,\operatorname{tw}}_{1,\{1\}}\mathcal{W}_{1}=-\frac{y^{3/2}-y^{-3/2}}{3i}\cdot\frac{3(y^{3/2}-y^{-3/2})}{i}\cdot\frac{-3}{(y^{1/2}-y^{-1/2})^{2}}=-3\left(\frac{y^{3/2}-y^{-3/2}}{y^{1/2}-y^{-1/2}}\right)^{2}.$
@@
$n_{0,4}=-192,\quad n_{1,4}=231,\quad n_{2,4}=-102,\quad n_{3,4}=15,$
@@
$\Omega_{4}^{\mathbb{P}^{2}}=\frac{1}{y^{3}}(1+y+4y^{2}+4y^{3}+4y^{4}+y^{5}+y^{6}){\frac{y^{6}-y^{-6}}{y^{\frac{1}{2}}-y^{-\frac{1}{2}}}}\,.$
@@
$\frac{-1}{(y^{\frac{1}{2}}-y^{-\frac{1}{2}})^{2}}\left(\frac{3d(-1)^{d^{2}+1}\Omega_{d}^{\mathbb{P}^{2}}\left(y^{\frac{1}{2}}\right)}{{\frac{\left(y^{\frac{1}{2}}\right)^{3d}-\left(y^{\frac{1}{2}}\right)^{-3d}}{y^{\frac{1}{2}}-y^{-\frac{1}{2}}}}}-\sum_{g\geq 0}n_{g,d}(-1)^{g}(y^{\frac{1}{2}}-y^{-\frac{1}{2}})^{2g}\right)=\sum_{T\in\mathcal{R}T_{d},\atop t^{-1}(0)\neq\emptyset}{\operatorname{Cont}}_{T}\,.$
@@
${\operatorname{Cont}}_{T_{1}}=-\frac{2\sin(3\hbar)}{6}F^{E,\operatorname{tw}}_{1,\{2\}}\mathcal{W}_{2}=-\frac{3(y^{3}-y^{-3})^{2}}{2(y-y^{-1})^{2}}+\frac{6(y^{3}-y^{-3})^{2}}{(y^{1/2}-y^{-1/2})^{2}}\,.$
@@
${\operatorname{Cont}}_{T_{2}}=-\frac{1}{2}\left(\frac{2\sin(3\hbar/2)}{3}\right)^{2}F^{E,\operatorname{tw}}_{1,\{1,1\}}(\mathcal{W}_{1})^{2}=\frac{9(y^{3/2}-y^{-3/2})^{4}}{2(y^{1/2}-y^{-1/2})^{4}}\,.$
@@
${\operatorname{Cont}}_{T_{3}}$ $=\left(y^{9/2}-y^{-9/2}\right)^{2}\left(-\frac{1}{(y^{3/2}-y^{-3/2})^{2}}-\frac{10y+10y^{-1}+7}{(y^{1/2}-y^{-1/2})^{2}}\right);$
@@
This method allows us to determine $\Omega_{d}^{\mathbb{P}^{2}}$ once the Gopakumar-Vafa invariants $n_{g,d^{\prime}}$ for $d^{\prime}\leq d$ are known. In the following table, we provide a list of $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ for $d\leq 10$, where $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ is given by (2). Actually, we have computed $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ for $d\leq 16$ using Gopakumar-Vafa invariants of local $\mathbb{P}^{2}$ in [HKR]. However, as the coefficients of $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ grow rapidly, we limit the list to $d\leq 10$ in this paper. A complete list of $\hat{\Omega}^{\mathbb{P}^{2}}_{d}$ for $d\leq 16$ is available at: